# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

### v1

@@ Acknowledgements.
- Arnaud (1996) Arnaud, K. A. 1996, in Astronomical Society of the Pacific Conference Series, Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes, 17
- Asahina et al. (2014) Asahina, Y., Ogawa, T., Kawashima, T., et al. 2014, ApJ, 789, 79
- Atapin & Fabrika (2016) Atapin, K. E. & Fabrika, S. N. 2016, AstL, 42, 517
- Begelman et al. (2006) Begelman, M. C., King, A. R., & Pringle, J. E. 2006, MNRAS, 370, 399
- Begelman et al. (1980) Begelman, M. C., Sarazin, C. L., Hatchett, S. P., McKee, C. F., & Arons, J. 1980, ApJ, 238, 722
- Blundell & Bowler (2004) Blundell, K. M. & Bowler, M. G. 2004, ApJL, 616, L159
- Blundell et al. (2011) Blundell, K. M., Schmidtobreick, L., & Trushkin, S. 2011, MNRAS, 417, 2401
- Borkowski et al. (2001) Borkowski, K. J., Lyerly, W. J., & Reynolds, S. P. 2001, ApJ, 548, 820
- Boumis et al. (2007) Boumis, P., Meaburn, J., Alikakos, J., et al. 2007, MNRAS, 381, 308
- Bregman et al. (2023) Bregman, J., Cen, R., Chen, Y., et al. 2023, SCPMA, 66, 299513
- Brinkmann et al. (1996) Brinkmann, W., Aschenbach, B., & Kawai, N. 1996, A&A, 312, 306
- Brinkmann et al. (1988) Brinkmann, W., Fink, H. H., Massaglia, S., Bodo, G., & Ferrari, A. 1988, A&A, 196, 313
- Brinkmann & Kawai (2000) Brinkmann, W. & Kawai, N. 2000, A&A, 363, 640
- Brinkmann et al. (1991) Brinkmann, W., Kawai, N., Matsuoka, M., & Fink, H. H. 1991, A&A, 241, 112
- Brinkmann et al. (2005) Brinkmann, W., Kotani, T., & Kawai, N. 2005, A&A, 431, 575
- Brinkmann et al. (2007) Brinkmann, W., Pratt, G. W., Rohr, S., Kawai, N., & Burwitz, V. 2007, A&A, 463, 611
- Broderick et al. (2018) Broderick, J. W., Fender, R. P., Miller-Jones, J. C. A., et al. 2018, MNRAS, 475, 5360
- Bykov et al. (2025) Bykov, A. M., Osipov, S. M., Romansky, V. I., et al. 2025, PhRvD, 112, 063017
- Cherepashchuk et al. (2023) Cherepashchuk, A., Belinski, A., Dodin, A., & Postnov, K. 2023, NewA, 103, 102060
- Cherepashchuk (1981) Cherepashchuk, A. M. 1981, MNRAS, 194, 761
- Cherepashchuk et al. (2025) Cherepashchuk, A. M., Dodin, A. V., & Postnov, K. A. 2025, arXiv, arXiv:2506.01106
- Cherepashchuk et al. (2018) Cherepashchuk, A. M., Esipov, V. F., Dodin, A. V., Davydov, V. V., & Belinskii, A. A. 2018, ARep, 62, 747
- Chi et al. (2024) Chi, Y.-H., Huang, J., Zhou, P., et al. 2024, ApJL, 975, L28
- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
- Churazov et al. (2021a) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021a, A&A, 651, A41
- Churazov et al. (2025) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Dolag, K. 2025, arXiv, arXiv:2507.19987
- Churazov et al. (2020) Churazov, E., Khabibullin, I., & Sunyaev, R. 2020, MNRAS, 495, L51
- Churazov et al. (2024) Churazov, E. M., Khabibullin, I. I., & Bykov, A. M. 2024, A&A, 688, A4
- Churazov et al. (2021b) Churazov, E. M., Khabibullin, I. I., Bykov, A. M., et al. 2021b, MNRAS, 507, 971
- Dubner et al. (1998) Dubner, G. M., Holdaway, M., Goss, W. M., & Mirabel, I. F. 1998, AJ, 116, 1842
- Eichler (1983) Eichler, D. 1983, ApJ, 272, 48
- Eikenberry et al. (2001) Eikenberry, S. S., Cameron, P. B., Fierce, B. W., et al. 2001, ApJ, 561, 1027
- Elston & Baum (1987) Elston, R. & Baum, S. 1987, AJ, 94, 1633
- Fabian & Rees (1979) Fabian, A. C. & Rees, M. J. 1979, MNRAS, 187, 13P
- Fabrika (2004) Fabrika, S. 2004, ASPRv, 12, 1
- Fabrika & Mescheryakov (2001) Fabrika, S. & Mescheryakov, A. 2001, in IAU Symposium, Vol. 205, Galaxies and their Constituents at the Highest Angular Resolutions, ed. R. T. Schilizzi, 268
- Farnes et al. (2017) Farnes, J. S., Gaensler, B. M., Purcell, C., et al. 2017, MNRAS, 467, 4777
- Filippova et al. (2006) Filippova, E., Revnivtsev, M., Fabrika, S., Postnov, K., & Seifina, E. 2006, A&A, 460, 125
- Fogantini et al. (2023) Fogantini, F. A., García, F., Combi, J. A., et al. 2023, A&A, 669, A149
- Forman et al. (1978) Forman, W., Jones, C., Cominsky, L., et al. 1978, ApJS, 38, 357
- Fragile et al. (2025) Fragile, P. C., Middleton, M. J., Bollimpalli, D. A., & Smith, Z. 2025, MNRAS, 540, 2820
- Fuchs et al. (2006) Fuchs, Y., Koch Miramond, L., & Ábrahám, P. 2006, A&A, 445, 1041
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Geldzahler et al. (1980) Geldzahler, B. J., Pauls, T., & Salter, C. J. 1980, A&A, 84, 237
- Goodall et al. (2011a) Goodall, P. T., Alouani-Bibi, F., & Blundell, K. M. 2011a, MNRAS, 414, 2838
- Goodall et al. (2011b) Goodall, P. T., Blundell, K. M., & Bell Burnell, S. J. 2011b, MNRAS, 414, 2828
- Goranskij (2011) Goranskij, V. 2011, PZ, 31, 5
- Green (2025) Green, D. A. 2025, JApA, 46, 14
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
- Gregory et al. (1996) Gregory, P. C., Scott, W. K., Douglas, K., & Condon, J. J. 1996, ApJS, 103, 427
- Grindlay et al. (1984) Grindlay, J. E., Band, D., Seward, F., et al. 1984, ApJ, 277, 286
- H. E. S. S. Collaboration et al. (2024) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2024, Sci, 383, 402
- Heinz & Sunyaev (2002) Heinz, S. & Sunyaev, R. 2002, A&A, 390, 751
- Hjellming & Johnston (1981a) Hjellming, R. M. & Johnston, K. J. 1981a, ApJL, 246, L141
- Hjellming & Johnston (1981b) Hjellming, R. M. & Johnston, K. J. 1981b, Nature, 290, 100
- Holden & Caswell (1969) Holden, D. J. & Caswell, J. L. 1969, MNRAS, 143, 407
- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
- Kaaret et al. (2017) Kaaret, P., Feng, H., & Roberts, T. P. 2017, ARA&A, 55, 303
- Kaaret et al. (2024) Kaaret, P., Ferrazzoli, R., Silvestri, S., et al. 2024, ApJL, 961, L12
- Katz (1980) Katz, J. I. 1980, ApJL, 236, L127
- Katz et al. (1982) Katz, J. I., Anderson, S. F., Margon, B., & Grandi, S. A. 1982, ApJ, 260, 780
- Kawai et al. (1989) Kawai, N., Matsuoka, M., Pan, H.-C., & Stewart, G. C. 1989, PASJ, 41, 491
- Khabibullin et al. (2022) Khabibullin, I., Churazov, E., & Sunyaev, R. 2022, MNRAS, 509, 6068
- Khabibullin et al. (2016) Khabibullin, I., Medvedev, P., & Sazonov, S. 2016, MNRAS, 455, 1414
- Khabibullin & Sazonov (2016) Khabibullin, I. & Sazonov, S. 2016, MNRAS, 457, 3963
- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
- Khabibullin et al. (2024) Khabibullin, I. I., Churazov, E. M., Chugai, N. N., et al. 2024, A&A, 689, A278
- Khabibullin & Sazonov (2012) Khabibullin, I. I. & Sazonov, S. Y. 2012, AstL, 38, 443
- Khabibullin & Sazonov (2017) Khabibullin, I. I. & Sazonov, S. Y. 2017, AstL, 43, 388
- Khabibullin & Sazonov (2019) Khabibullin, I. I. & Sazonov, S. Y. 2019, AstL, 45, 282
- Kimura et al. (2020) Kimura, S. S., Murase, K., & Mészáros, P. 2020, ApJ, 904, 188
- Kirshner & Chevalier (1980) Kirshner, R. P. & Chevalier, R. A. 1980, ApJL, 242, L77
- Konigl (1983) Konigl, A. 1983, MNRAS, 205, 471
- Kotani et al. (1996) Kotani, T., Kawai, N., Matsuoka, M., & Brinkmann, W. 1996, PASJ, 48, 619
- Kraft et al. (2022) Kraft, R., Markevitch, M., Kilbourne, C., et al. 2022, arXiv, arXiv:2211.09827
- Kubota et al. (2010) Kubota, K., Ueda, Y., Kawai, N., et al. 2010, PASJ, 62, 323
- Leibowitz (1984) Leibowitz, E. M. 1984, MNRAS, 210, 279
- LHAASO Collaboration (2024) LHAASO Collaboration. 2024, arXiv, arXiv:2410.08988
- Lopez et al. (2006) Lopez, L. A., Marshall, H. L., Canizares, C. R., Schulz, N. S., & Kane, J. F. 2006, ApJ, 650, 338
- Margon (1984) Margon, B. 1984, ARA&A, 22, 507
- Margon et al. (1979) Margon, B., Ford, H. C., Katz, J. I., et al. 1979, ApJL, 230, L41
- Marshall et al. (2013) Marshall, H. L., Canizares, C. R., Hillwig, T., et al. 2013, ApJ, 775, 75
- Marshall et al. (2002) Marshall, H. L., Canizares, C. R., & Schulz, N. S. 2002, ApJ, 564, 941
- Medvedev & Fabrika (2010) Medvedev, A. & Fabrika, S. 2010, MNRAS, 402, 479
- Medvedev et al. (2019) Medvedev, P. S., Khabibullin, I. I., & Sazonov, S. Y. 2019, AstL, 45, 299
- Medvedev et al. (2018) Medvedev, P. S., Khabibullin, I. I., Sazonov, S. Y., Churazov, E. M., & Tsygankov, S. S. 2018, AstL, 44, 390
- Medvedev et al. (2022) Medvedev, P. S., Khabibullin, I. I., Semena, A. N., et al. 2022, AstL, 48, 389
- Middleton et al. (2021) Middleton, M. J., Walton, D. J., Alston, W., et al. 2021, MNRAS, 506, 1045
- Migliari et al. (2005) Migliari, S., Fender, R. P., Blundell, K. M., Méndez, M., & van der Klis, M. 2005, MNRAS, 358, 860
- Miller-Jones et al. (2008) Miller-Jones, J. C. A., Migliari, S., Fender, R. P., et al. 2008, ApJ, 682, 1141
- Monceau-Baroux et al. (2015) Monceau-Baroux, R., Porth, O., Meliani, Z., & Keppens, R. 2015, A&A, 574, A143
- Namiki et al. (2000) Namiki, M., Kawai, N., Kotani, T., Mamauchi, S., & Brinkmann, W. 2000, AdSpR, 25, 709
- Ohmura et al. (2021) Ohmura, T., Ono, K., Sakemi, H., et al. 2021, ApJ, 910, 149
- Ohsuga & Mineshige (2011) Ohsuga, K. & Mineshige, S. 2011, ApJ, 736, 2
- Panferov (2014) Panferov, A. 2014, A&A, 562, A130
- Panferov (2017) Panferov, A. A. 2017, A&A, 599, A77
- Panferov & Kabanov (2013) Panferov, A. A. & Kabanov, A. A. 2013, arXiv, arXiv:1306.4486
- Peretti et al. (2025) Peretti, E., Petropoulou, M., Vasilopoulos, G., & Gabici, S. 2025, A&A, 698, A188
- Peter & Eichler (1993) Peter, W. & Eichler, D. 1993, ApJ, 417, 170
- Poutanen et al. (2007) Poutanen, J., Lipunova, G., Fabrika, S., Butkevich, A. G., & Abolmasov, P. 2007, MNRAS, 377, 1187
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Predehl et al. (2020) Predehl, P., Sunyaev, R. A., Becker, W., et al. 2020, Nature, 588, 227
- Rosado et al. (2021) Rosado, M., Sánchez-Cruces, M., Ambrocio-Cruz, P., & Trejo, D. 2021, MNRAS, 506, 4263
- Safi-Harb et al. (2022) Safi-Harb, S., Mac Intyre, B., Zhang, S., et al. 2022, ApJ, 935, 163
- Safi-Harb & Ögelman (1997) Safi-Harb, S. & Ögelman, H. 1997, ApJ, 483, 868
- Safi-Harb & Petre (1999) Safi-Harb, S. & Petre, R. 1999, ApJ, 512, 784
- Sakai et al. (2025) Sakai, Y., Yamada, S., Sakemi, H., et al. 2025, arXiv, arXiv:2507.19042
- Sakemi et al. (2021) Sakemi, H., Omae, R., Ohmura, T., & Machida, M. 2021, PASJ, 73, 530
- Sarazin et al. (1980) Sarazin, C. L., Begelman, M. C., & Hatchett, S. P. 1980, ApJL, 238, L129
- Seward et al. (1980) Seward, F., Grindlay, J., Seaquist, E., & Gilmore, W. 1980, Nature, 287, 806
- Seward et al. (1976) Seward, F. D., Page, C. G., Turner, M. J. L., & Pounds, K. A. 1976, MNRAS, 175, 39P
[…]

### v2

@@ Acknowledgements.
- XSPEC: The First Ten Years. In Astronomical Data Analysis Software and Systems V, G. H. Jacoby and J. Barnes (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 101, pp. 17. External Links: ADS entry Cited by: Appendix B.
- Magnetohydrodynamic Simulations of a Jet Drilling an H I Cloud: Shock Induced Formation of Molecular Clouds and Jet Breakup. ApJ 789 (1), pp. 79. External Links: Document, 1407.2381, ADS entry Cited by: §6.
- X-ray variability of SS 433: Evidence for supercritical accretion. AstL 42 (8), pp. 517–530. External Links: Document, 1607.04185, ADS entry Cited by: §1.
- The nature of SS433 and the ultraluminous X-ray sources. MNRAS 370 (1), pp. 399–404. External Links: Document, astro-ph/0604497, ADS entry Cited by: §1, footnote 2.
- Beam models for SS 433.. ApJ 238, pp. 722–730. External Links: Document, ADS entry Cited by: §1, §6.
- Symmetry in the Changing Jets of SS 433 and Its True Distance from Us. ApJL 616 (2), pp. L159–L162. External Links: Document, astro-ph/0410456, ADS entry Cited by: §1.
- SS433’s accretion disc, wind and jets: before, during and after a major flare. MNRAS 417 (4), pp. 2401–2410. External Links: Document, 1104.2917, ADS entry Cited by: §1.
- Supernova Remnants in the Sedov Expansion Phase: Thermal X-Ray Emission. ApJ 548 (2), pp. 820–835. External Links: Document, astro-ph/0008066, ADS entry Cited by: Appendix B.
- Deep optical observations of the interaction of the SS 433 microquasar jet with the W50 radio continuum shell. MNRAS 381 (1), pp. 308–318. External Links: Document, 0707.4243, ADS entry Cited by: §1, §3, §6.
- Scientific objectives of the Hot Universe Baryon Surveyor (HUBS) mission. SCPMA 66 (9), pp. 299513. External Links: Document, 2307.05672, ADS entry Cited by: §6.
- ROSAT observations of the W 50/SS 433 system.. A&A 312, pp. 306–316. External Links: ADS entry Cited by: §1, §1, §2, §2, §3, §5.2.3.
- The thermal evolution of the jets in SS 433.. A&A 196, pp. 313–326. External Links: ADS entry Cited by: Appendix A, §1, §5.1.
- The X-ray emission of SS 433.. A&A 241, pp. 112. External Links: ADS entry Cited by: §5.1.
- The jets of SS 433: second order effects. A&A 363, pp. 640–646. External Links: ADS entry Cited by: Appendix A.
- XMM-Newton observations of SS 433 I. EPIC spectral analysis. A&A 431, pp. 575–586. External Links: Document, ADS entry Cited by: §1, §1, §2, Figure 6, §5.1.
- XMM-Newton observations of the eastern jet of SS 433. A&A 463 (2), pp. 611–619. External Links: Document, astro-ph/0610781, ADS entry Cited by: §1, §3, §4, §5.2.4, §6.
- LOFAR 150-MHz observations of SS 433 and W 50. MNRAS 475 (4), pp. 5360–5377. External Links: Document, 1802.03406, ADS entry Cited by: §1.
- PeV particle acceleration and nonthermal emission in the minimalist model of the extended jets in W50/SS433. PhRvD 112 (6), pp. 063017. External Links: Document, 2509.09883, ADS entry Cited by: §1, §6.
- Spectroscopic Monitoring of SS 433. Stability of Parameters of the Kinematic Model over 40 Years. ARep 62 (11), pp. 747–763. External Links: Document, ADS entry Cited by: §1.
- SS 433 as an Eclipsing Binary.. MNRAS 194, pp. 761–769. External Links: Document, ADS entry Cited by: §1.
- Evolutionary increase of the orbital separation and change of the Roche lobe size in SS433. NewA 103, pp. 102060. External Links: Document, 2305.07093, ADS entry Cited by: §1.
- Unique microquasar SS433: new results, new issues. PhyU 68 (10), pp. 1042–1060. External Links: Document, 2506.01106, ADS entry Cited by: §1.
- An X-Ray Shell Reveals the Supernova Explosion for Galactic Microquasar SS 433. ApJL 975 (2), pp. L28. External Links: Document, 2410.06510, ADS entry Cited by: §1, §1, §2, §5.2.3.
- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA. II. Shock and relic. A&A 670, pp. A156. External Links: Document, 2205.07511, ADS entry Cited by: §2.
- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA. I. X-ray morphology, recent merger, and radio halo connection. A&A 651, pp. A41. External Links: Document, 2012.11627, ADS entry Cited by: §2.
- Does Nature use neutral beams for interstellar plasma heating around compact objects?. MNRAS 495 (1), pp. L51–L55. External Links: Document, 2002.02027, ADS entry Cited by: §1, §6.
- SRG/eROSITA discovery of a large circular SNR candidate G116.6-26.1: SN Ia explosion probing the gas of the Milky Way halo?. MNRAS 507 (1), pp. 971–982. External Links: Document, 2106.09454, ADS entry Cited by: §2.
- Minimalist model of the W50/SS433 extended X-ray jet: Anisotropic wind with recollimation shocks. A&A 688, pp. A4. External Links: Document, 2401.14770, ADS entry Cited by: §1, §2, §3, Figure 8, §6.
- Perseus cluster in its X-ray entirety with SRG/eROSITA et al. I. Merger and Radio-Uroboroses. arXiv, pp. arXiv:2507.19987. External Links: Document, 2507.19987, ADS entry Cited by: §2.
- A High-Resolution Radio Study of the W50-SS 433 System and the Surrounding Medium. AJ 116 (4), pp. 1842–1855. External Links: Document, ADS entry Cited by: §1, Figure 8, §6.
- Focusing of high-mach number jets by an ambient medium.. ApJ 272, pp. 48–53. External Links: Document, ADS entry Cited by: §6, §6.
- Twenty Years of Timing SS 433. ApJ 561 (2), pp. 1027–1033. External Links: Document, astro-ph/0107296, ADS entry Cited by: §2.
- High-Resolution Radio Observations of W50, the Remnant Associated with SS 433. AJ 94, pp. 1633. External Links: Document, ADS entry Cited by: §1.
- SS 433: a double jet in action?. MNRAS 187, pp. 13P–16. External Links: Document, ADS entry Cited by: §1.
- Face-on SS 433 stars as a possible new type of extragalactic X-ray sources. In Galaxies and their Constituents at the Highest Angular Resolutions, R. T. Schilizzi (Ed.), IAU Symposium, Vol. 205, pp. 268. External Links: Document, astro-ph/0103070, ADS entry Cited by: §1, footnote 2.
- The jets and supercritical accretion disk in SS433. ASPRv 12, pp. 1–152. External Links: Document, astro-ph/0603390, ADS entry Cited by: §1.
- Interacting large-scale magnetic fields and ionized gas in the W50/SS433 system. MNRAS 467 (4), pp. 4777–4801. External Links: Document, 1604.06552, ADS entry Cited by: §1, §3, §6.
- Diagnostics of SS433 with the RXTE. A&A 460 (1), pp. 125–131. External Links: Document, astro-ph/0609367, ADS entry Cited by: §1.
- A NuSTAR view of SS433. Precessional evolution of the jet-disk system. A&A 669, pp. A149. External Links: Document, 2210.09390, ADS entry Cited by: §1, footnote 2.
- The fourth Uhuru catalog of X-ray sources.. ApJS 38, pp. 357–412. External Links: Document, ADS entry Cited by: §1.
- Long time-scale numerical simulations of large supercritical accretion discs. MNRAS 540 (3), pp. 2820–2829. External Links: Document, 2505.08859, ADS entry Cited by: §1.
- SS 433: a phenomenon imitating a Wolf-Rayet star. A&A 445 (3), pp. 1041–1052. External Links: Document, astro-ph/0512190, ADS entry Cited by: §1.
- Gaia Data Release 3. Summary of the content and survey properties. A&A 674, pp. A1. External Links: Document, 2208.00211, ADS entry Cited by: footnote 1.
- Continuum observations of the SNR W50 and G 74.9+1.2 at 2695 MHz.. A&A 84, pp. 237–244. External Links: ADS entry Cited by: §1.
- When microquasar jets and supernova collide: hydrodynamically simulating the SS 433-W 50 interaction. MNRAS 414 (4), pp. 2838–2859. External Links: Document, 1101.3486, ADS entry Cited by: §1, §1, §2, §4, §6, §6.
- Probing the history of SS 433’s jet kinematics via decade-resolution radio observations of W 50. MNRAS 414 (4), pp. 2828–2837. External Links: Document, 1103.5658, ADS entry Cited by: §1.
- Photometric Mass Estimate for the Compact Component of SS 433: And Yet It Is a Neutron Star. PZ 31 (5), pp. 5. External Links: Document, 1110.5304, ADS entry Cited by: §1.
- A colour scheme for the display of astronomical intensity images. BASI 39, pp. 289–295. External Links: Document, 1108.5083, ADS entry Cited by: Data availability.
- An updated catalogue of 310 Galactic supernova remnants and their statistical properties. JApA 46 (1), pp. 14. External Links: Document, 2411.03367, ADS entry Cited by: Figure 1.
- A 3D Dust Map Based on Gaia, Pan-STARRS 1, and 2MASS. ApJ 887 (1), pp. 93. External Links: Document, 1905.02734, ADS entry Cited by: Figure 9, Appendix C, §2.
- The GB6 Catalog of Radio Sources. ApJS 103, pp. 427. External Links: Document, ADS entry Cited by: Figure 1, §3, Figure 2.
- The central X-ray source in SS 433.. ApJ 277, pp. 286–295. External Links: Document, ADS entry Cited by: §1.
- Acceleration and transport of relativistic electrons in the jets of the microquasar SS 433. Sci 383 (6681), pp. 402–406. External Links: Document, 2401.16019, ADS entry Cited by: §1, §5.2.4.
- Cosmic rays from microquasars: A narrow component to the CR spectrum?. A&A 390, pp. 751–766. External Links: Document, astro-ph/0204183, ADS entry Cited by: §1.
- An analysis of the proper motions of SS 433 radio jets.. ApJL 246, pp. L141–L145. External Links: Document, ADS entry Cited by: §1.
- Structure, strength and polarization changes in radio source SS 433.. Nature 290, pp. 100–107. External Links: Document, ADS entry Cited by: §1.
- 178 MHz observations of galactic radio sources.. MNRAS 143, pp. 407–435. External Links: Document, ADS entry Cited by: §1.
- Super-Eddington Accretion Disks around Supermassive Black Holes. ApJ 880 (2), pp. 67. External Links: Document, 1709.02845, ADS entry Cited by: §1.
- Ultraluminous X-Ray Sources. ARA&A 55 (1), pp. 303–341. External Links: Document, 1703.10728, ADS entry Cited by: §1.
- X-Ray Polarization of the Eastern Lobe of SS 433. ApJL 961 (1), pp. L12. External Links: Document, 2311.16313, ADS entry Cited by: §1, §1, §3, §5.2.4, §6.
- Nodding motions of accretion rings and disks : a short-term period inSS 433.. ApJ 260, pp. 780–793. External Links: Document, ADS entry Cited by: §1.
- Acceleration, radiation and precession in SS 433.. ApJL 236, pp. L127–L130. External Links: Document, ADS entry Cited by: §1.
- GINGA observations of the X-ray eclipse of SS 433.. PASJ 41, pp. 491–507. External Links: ADS entry Cited by: §1, §2, §5.1.
- SRG/eROSITA discovery of a radio-faint X-ray candidate supernova remnant SRGe J003602.3+605421 = G121.1-1.9. MNRAS 521 (4), pp. 5536–5556. External Links: Document, 2207.00064, ADS entry Cited by: §2.
- X-ray line formation in the spectrum of SS 433. AstL 38 (7), pp. 443–467. External Links: Document, 1206.0451, ADS entry Cited by: Appendix A.
- Identification of X-ray lines in the spectrum of the arcsec-scale precessing jets of SS 433. AstL 43 (6), pp. 388–399. External Links: Document, 1701.05884, ADS entry Cited by: §1, §4.
- Constraints on the Collimated X-ray Emission of SS 433 from the Reflection on Molecular Clouds. AstL 45 (5), pp. 282–298. External Links: Document, 1905.09667, ADS entry Cited by: §1, §5.2.4.
- Thermal X-ray emission from a baryonic jet: a self-consistent multicolour spectral model. MNRAS 455 (2), pp. 1414–1427. External Links: Document, 1510.05563, ADS entry Cited by: Appendix A, §1, Figure 6, §5.1.
- Is SS 433 a misaligned ultraluminous X-ray source? Constraints from its reflected signal in the Galactic plane. MNRAS 457 (4), pp. 3963–3974. External Links: Document, 1512.02616, ADS entry Cited by: §1, §5.2.4, footnote 2.
- SRG/eROSITA view of X-ray reflection in the Central Molecular Zone: a snapshot in September-October 2019. MNRAS 509 (4), pp. 6068–6076. External Links: Document, 2111.07899, ADS entry Cited by: §2.
- Study of X-ray emission from the S147 nebula by SRG/eROSITA: Supernova-in-the-cavity scenario. A&A 689, pp. A278. External Links: Document, 2401.17261, ADS entry Cited by: §2.
- Deciphering the Origin of the GeV-TeV Gamma-Ray Emission from SS 433. ApJ 904 (2), pp. 188. External Links: Document, 2008.04515, ADS entry Cited by: §1.
- Spectrophotometry of W 50 : the SNR around SS 433.. ApJL 242, pp. L77–L81. External Links: Document, ADS entry Cited by: §1.
- W 50 - a stellar-wind bubble in a three-phase intestellar medium ?. MNRAS 205, pp. 471–485. External Links: Document, ADS entry Cited by: §6.
- Iron-Line Diagnostics of the Jets of SS 433. PASJ 48, pp. 619–629. External Links: Document, ADS entry Cited by: §1, §5.1.
- Line Emission Mapper (LEM): Probing the physics of cosmic ecosystems. arXiv, pp. arXiv:2211.09827. External Links: Document, 2211.09827, ADS entry Cited by: §6.
- Suzaku X-Ray and Optical Spectroscopic Observations of SS 433 in the 2006 April Multiwavelength Campaign. PASJ 62, pp. 323. External Links: Document, 1001.2989, ADS entry Cited by: §1.
- A geometrical model for the SS 433 system.. MNRAS 210, pp. 279–301. External Links: Document, ADS entry Cited by: §1.
- Ultrahigh-Energy Gamma-ray Emission Associated with Black Hole-Jet Systems. National Science Review 12 (12), pp. nwaf496. External Links: Document, 2410.08988, ADS entry Cited by: §1.
- Determining the Nature of the SS 433 Binary from an X-Ray Spectrum during Eclipse. ApJ 650 (1), pp. 338–349. External Links: Document, astro-ph/0605574, ADS entry Cited by: §1, §2, §5.1.
- The bizarre spectrum of SS 433.. ApJL 230, pp. L41–L45. External Links: Document, ADS entry Cited by: §1.
- Observations of SS 433. ARA&A 22, pp. 507–536. External Links: Document, ADS entry Cited by: §1.
- Multiwavelength Observations of the SS 433 Jets. ApJ 775 (1), pp. 75. External Links: Document, 1307.8427, ADS entry Cited by: §1, §1, §2, §5.1.
- The High-Resolution X-Ray Spectrum of SS 433 Using the Chandra HETGS. ApJ 564 (2), pp. 941–952. External Links: Document, astro-ph/0108206, ADS entry Cited by: Appendix A, §1, §4, Figure 6, §5.1.
- Evidence of supercritical disc funnel radiation in X-ray spectra of SS 433. MNRAS 402 (1), pp. 479–491. External Links: Document, 0910.3186, ADS entry Cited by: Appendix A, §1, Figure 6, §5.1, footnote 2.
- An Upper Limit on Nickel Overabundance in the Supercritical Accretion Disk Wind of SS 433 from X-ray Spectroscopy. AstL 44 (6), pp. 390–410. External Links: Document, 1804.01828, ADS entry Cited by: Appendix A, §1.
- Diagnostics of Parameters for the X-ray Jets of SS 433 from High-Resolution Chandra Spectroscopy. AstL 45 (5), pp. 299–320. External Links: Document, 2005.12416, ADS entry Cited by: Appendix A, §1, §2, Figure 6, §5.1.
- The X-ray Jets of SS 433 in the Period of Flaring Activity in the Summer of 2018. AstL 48 (7), pp. 389–405. External Links: Document, ADS entry Cited by: §1.
- NuSTAR reveals the hidden nature of SS433. MNRAS 506 (1), pp. 1045–1058. External Links: Document, 1810.10518, ADS entry Cited by: §1, footnote 2.
- Rapid variability of the arcsec-scale X-ray jets of SS 433. MNRAS 358 (3), pp. 860–868. External Links: Document, astro-ph/0501097, ADS entry Cited by: §1.
- Coupled Radio and X-Ray Emission and Evidence for Discrete Ejecta in the Jets of SS 433. ApJ 682 (2), pp. 1141–1151. External Links: Document, 0804.1337, ADS entry Cited by: §1.
- The SS433 jet from subparsec to parsec scales. A&A 574, pp. A143. External Links: Document, ADS entry Cited by: §6, §6.
- X-Ray Lobes of W50/SS433 System. AdSpR 25 (3-4), pp. 709–712. External Links: Document, ADS entry Cited by: §1, §3.
- Continuous Jets and Backflow Models for the Formation of W50/SS 433 in Magnetohydrodynamics Simulations. ApJ 910 (2), pp. 149. External Links: Document, 2102.06728, ADS entry Cited by: §6.
- Global Structure of Three Distinct Accretion Flows and Outflows around Black Holes from Two-dimensional Radiation-magnetohydrodynamic Simulations. ApJ 736 (1), pp. 2. External Links: Document, 1105.5474, ADS entry Cited by: §1.
- Notes on non-thermal X-ray radiation of radio supernova remnant W50 and collimated radiation of SS433. arXiv, pp. arXiv:1306.4486. External Links: Document, 1306.4486, ADS entry Cited by: §5.2.4.
- Simulation of kinematics of SS 433 radio jets that interact with the ambient medium. A&A 562, pp. A130. External Links: Document, 1306.3599, ADS entry Cited by: §1, §6.
- Jets of SS 433 on scales of dozens of parsecs. A&A 599, pp. A77. External Links: Document, 1607.02043, ADS entry Cited by: §1, §6, §6.
- Particle acceleration and multi-messenger radiation from ultra-luminous X-ray sources: A new class of Galactic PeVatrons. A&A 698, pp. A188. External Links: Document, 2411.08762, ADS entry Cited by: §1.
- Hydrodynamic Collimation of Precessing Jets. ApJ 417, pp. 170. External Links: Document, ADS entry Cited by: §6, §6.
- Supercritically accreting stellar mass black holes as ultraluminous X-ray sources. MNRAS 377 (3), pp. 1187–1194. External Links: Document, astro-ph/0609274, ADS entry Cited by: §1, footnote 2.
- The eROSITA X-ray telescope on SRG. A&A 647, pp. A1. External Links: Document, 2010.03477, ADS entry Cited by: §1, §2, §2.
- Detection of large-scale X-ray bubbles in the Milky Way halo. Nature 588 (7837), pp. 227–231. External Links: Document, 2012.05840, ADS entry Cited by: §2.
- Kinematics of the western filament of W50. MNRAS 506 (3), pp. 4263–4274. External Links: Document, ADS entry Cited by: §1.
- Hard X-Ray Emission from the Eastern Jet of SS 433 Powering the W50 “Manatee” Nebula: Evidence for Particle Reacceleration. ApJ 935 (2), pp. 163. External Links: Document, 2207.00573, ADS entry Cited by: §1, §1, §1, §2, §4, §5.2.4, §6.
- ROSAT and ASCA Observations of W50 Associated with the Peculiar Source SS 433. ApJ 483 (2), pp. 868–881. External Links: Document, ADS entry Cited by: §1, §3.
- Rossi X-Ray Timing Explorer Observations of the Eastern Lobe of W50 Associated with SS 433. ApJ 512 (2), pp. 784–792. External Links: Document, ADS entry Cited by: §1, §3.
- Arcsecond-scale X-ray imaging and spectroscopy of SS 433 with the Chandra High-Energy Transmission Grating. PASJ 77 (5), pp. 1113–1125. External Links: Document, 2507.19042, ADS entry Cited by: §1.
- Energy estimation of high-energy particles associated with the SS 433/W 50 system through radio observation at 1.4 GHz. PASJ 73 (3), pp. 530–544. External Links: Document, 2103.05578, ADS entry Cited by: §1.
[…]